show all work. say a researcher is interested in testing the effectiveness of a new stress reduction method. She thinks, or predicts based on previous research, that the new method should reduce stress. one randonmly selected person (our subject) is trained in this method and his stress is measured (low values mean low stress). It is known that stress level of people in the general population has a mean of 40 and a standard deviation of 10. the score of the subject who received the method was X= 25.

1) complete the steps of hypothesis testing using the p<.05 significance level (Alpha). show each step) Given: the cutoff critical Z value that cuts off the 5% tail of the distribution =-1.64.

1a) What do you conclude about the anxiety reducing effectiveness of the method?

1b) Did you make a directional or non directional hypothesis? why?

1c) did you do a one-tailed or two-tailed test? why?
Answer #1

Lets put out the hypothesis:

H0: Mu < 40

Ha: Mu >= 40

1. First calc. the te statistic = X-Mu / Sigma = (25-40)/10 = -1.5

Now, since our test statistic is more than -1.64 we have the fact that the test statistic is not in the critical range

Hence, we fail to reject Null hypothesis i.e. Pop, mean = 40

2. We conclude that the anxiety reducing effectiveness method has no significant effect on anxeity, i.e. the anxeity continous to be the same even after this treatment

3. It is a directional / 1 tailed hypothesis

4. I did a 1 tailed test as I wanted to check wether my anxeity is lesser than the pop. anxeity of 40. So, there was only 1 tailed involved.